The ratio of men to women working at a factory Smythe Manufacturing Plant is 4:5. If there are 36 men, how many women are there?
Answer by ewatrrr(24785)

 
Hi
Let x represent the number of women
Note: ratios Set Up Proportions!
4/5 = 36/x | Cross Multiplying
4x = 5*36
x = 5*36/4
x = 45
CHECKING our Answer***
4/5 = 36/45